OpenStudy (anonymous):

needng help wth homework, I have the numbers 20, 10, 3, 2, and 1 and they need to equal 100% and need to become percents but I can never understand said concept. Could anyone help me?

OpenStudy (jango_in_dtown):

@Mokiandmidnight may I help?

OpenStudy (anonymous):

shoot for it, I tried and i just converted them into percents and they equaled to 36%

OpenStudy (jango_in_dtown):

see sum of the numbers =100% now sum of numers =36 so 36=100% therefore the number 20=(20/36*100 )%=(55 5/9 percents) 10=(10/36*100)%= (27 7/9) percents 3=(3/36*100)%=(8 1/3) percents 2=(2/36*100)=(5 5/9) percents 1=(1/36*100) %=(2 7/9) percents

OpenStudy (anonymous):

So anotherwords from what im gathering 20 is 55 5/9%, 10 is 27 7/9%, 3 is 8 1/3% and so on?

OpenStudy (anonymous):

Or should those numbers be rounded up?

OpenStudy (jango_in_dtown):

correct... I kept the solution in mixed fractions. You may keep them as it is or make them into partial fractions but cannot ronudup.. or else they wont add up to 100%

OpenStudy (anonymous):

Alright Thank you!